OpenStudy (anonymous):

PLEASE Which set of numbers could represent the lengths of the sides of a right triangle? 4,√65 , 9 6,√82 , 13 5,√55 , 10 11,√47 , 13

OpenStudy (anonymous):

c2=a2+b2

OpenStudy (anonymous):

still dont get it

OpenStudy (anonymous):

9^2=4^2+65

OpenStudy (anonymous):

answr is a